# Towards Federated Graph Learning in One-shot Communication

### v5

@@ Abstract
Federated Graph Learning (FGL) has emerged as a promising paradigm for breaking data silos among distributed private graphs. In practical scenarios involving heterogeneous distributed graph data, personalized Federated Graph Learning (pFGL) aims to enhance model utility by training personalized models tailored to client needs. However, existing pFGL methods often require numerous communication rounds under heterogeneous graphs, leading to significant communication overhead and security concerns. While One-shot Federated Learning (OFL) enables collaboration in a single round, existing OFL methods are designed for image-centric tasks and ineffective for graph data, leaving a critical gap in the field. Additionally, personalized models derived from existing methods suffer from bias, failing to effectively generalize to the minority. To address these challenges, we propose the first One-shot personalized Federated Graph Learning method (O-pFGL) for node classification, compatible with Secure Aggregation protocols for privacy preservation. Specifically, for effective graph learning in one communication round, our method estimates and aggregates class-wise feature distribution statistics to construct a global pseudo-graph on the server, facilitating the training of a global graph model. To mitigate bias, we introduce a two-stage personalized training approach that adaptively balances local personal information and global insights from the pseudo-graph, improving both personalization and generalization. Extensive experiments on 12 multi-scale graph datasets demonstrate that our method significantly outperforms state-of-the-art baselines across various settings.
Machine Learning, ICML
@@ 1 Introduction
Graphs are widely employed to model complex relationships between entities across a variety of domains (He et al., 2023; Hyun et al., 2023; Bang et al., 2023). While various algorithms and models (Gasteiger et al., 2018; Iscen et al., 2019; Wu et al., 2020) have been proposed, most of them assume that graph data from different sources are managed centrally. However, collecting and managing such graph data is often costly, impractical, and poses significant privacy risks in many sensitive scenarios (Fu et al., 2022).
Figure 1: The key limitations of existing methods. Our method stands out as the only method capable of supporting one-shot federated learning for graph data while accommodating model heterogeneity and Secure Aggregation.
@@
Figure 2: (a) A toy example explains the major classes and minor classes. (b) The empirical analysis of a client on the CiteSeer dataset under the Louvain partition to illustrate L3. Left: Existing methods result in a biased model that neglects the minority. Middle: The performance of existing and our methods in class 1 with 50 samples. Right: The performance of existing and our methods in class 2 with 14 samples. Our method effectively mitigates the performance gap.
@@
To address the tension between the need for vast datasets and the growing demand for privacy protection, Federated Graph Learning (FGL) has emerged as a promising solution, which enables collaborative training on distributed graph data without requiring centralized management (Pan et al., 2023; Gu et al., 2023). While FGL shows its potential, it faces two major limitations: (1) Typical FGL struggles to effectively personalize models for joint improvement across heterogeneous clients’ graph data. This heterogeneity arises from variations of both node and structural properties, with adjacent nodes influencing each other, further complicating personalization. (2) FGL requires iterative communication between clients and the server, raising efficiency concerns due to communication overhead (Kairouz et al., 2021; Dai et al., 2022) and security concerns (e.g. man-in-the-middle attacks (Wang et al., 2020; Mothukuri et al., 2021; Zhu et al., 2024a)). These limitations can be mitigated by integrating personalized federated graph learning (pFGL) and one-shot federated learning (OFL), which offer novel solutions to improve communication efficiency and model personalization.
@@
To address these challenges, we explore the integration of pFGL and OFL. The pFGL methods allow for personalized model training by focusing on individual client graph data. OFL methods limit communication to a single round, thereby reducing the communication overhead and circumventing security risks such as man-in-the-middle attacks (Guha et al., 2019; Zhang et al., 2022). However, both existing pFGL and OFL methods face challenges in FGL. Specifically, they exhibit the following limitations:
@@
L1: Ineffectiveness for graph learning in one communication round. Existing pFGL methods (Baek et al., 2023; Li et al., 2024b; Zhu et al., 2024b) rely on iterative communication for optimization, making them ineffective within one communication round. Most methods also fail to support heterogeneous models, as they primarily rely on aggregating model parameters. Meanwhile, existing OFL methods, which focus primarily on image-centric tasks, are incompatible with graph data. The ensemble-based OFL methods (Zhang et al., 2022; Dai et al., 2024) can invert images from ensemble models, but struggle with graph data due to the inter-dependencies between nodes. The distillation-based OFL faces the problems of distilling inter-related graphs. The generative-based OFL methods (Heinbaugh et al., 2023; Yang et al., 2024b, a) rely on pre-trained generative models (e.g. Stable Diffusion (Rombach et al., 2022)), which are not suitable for graph generation.
@@
L2: Incompatibility with Secure Aggregation. Secure Aggregation is a widely used protocol (Bonawitz et al., 2017; So et al., 2022) designed to safeguard client privacy during the federated aggregation process. However, most pFGL and OFL methods require clients to upload raw model parameters or auxiliary information independently, which complicates the aggregation process (e.g. ensemble or similarity calculation) and hinders the implementation of privacy-preserving weighted averages.
@@
L3: Imbalance between personalization and local generalization in personalization. Local graph data often exhibit a significant imbalance in both quantity and topology, as illustrated in Figure 2a. Personalized models tend to be biased toward major classes (with larger quantities and high homophily to form compact communities), while neglecting minor classes with fewer samples and often surrounded by nodes from major classes. This imbalance leads to a degradation in overall model performance, particularly in underrepresented classes. A detailed empirical analysis is provided in Figure 2b, which shows that existing personalized methods result in heavy bias toward majority classes (class 3 and class 5, with a total of 180 samples) while failing to adequately represent minority classes (class 1 and class 2, with only 64 samples), highlighting the limitations of current methods in handling imbalanced graph data.
@@
To overcome these limitations, we propose a One-shot personalized Federated Graph Learning method (O-pFGL). Our method supports model heterogeneity and Secure Aggregation protocols for privacy preservation. Specifically, for L1 and L2, clients upload estimated feature distribution statistics, augmented by our proposed homophily-guided reliable node expansion strategy, instead of raw model parameters. The server securely aggregates these statistics to generate a global pseudo-graph, which is subsequently utilized to train personalized models on clients, eliminating the need for direct model parameter exchanges and preventing model theft. To address L3, we propose a two-stage personalized training approach. In the first stage, clients train a global model based on the pseudo-graph, which serves as both a fixed teacher capturing global knowledge and a start-point for subsequent fine-tuning. In the second stage, clients fine-tune the global model on local graph data with the node-adaptive distillation, which leverages the global knowledge from the teacher model to counteract biases toward majority, ensuring effective learning for minority and achieving a robust balance between personalization and generalization.
@@
Our method significantly outperforms state-of-the-art baselines through comprehensive experiments across various settings, including different graph scales, partition strategies, learning paradigms, and evaluation metrics. Experimental results demonstrate the superiority of our method in achieving a one-shot personalized federated graph learning solution that balances efficiency, privacy, and personalization. To sum up, our contributions are listed as follows:
@@ 2.1 Federated Graph Learning
From the graph level (Xie et al., 2021; Tan et al., 2023), each client possesses multiple independent graphs (e.g. molecular graphs) and target graph-level downstream tasks. From the subgraph level, each client possesses a subgraph of an implicit global graph. FedPUB (Baek et al., 2023) and FedGTA (Li et al., 2024b) personalizes model by similarity of functional embedding and mixed moments of neighbor features. AdaFGL (Li et al., 2024a) studies the structure non-IID problem. FedTAD (Zhu et al., 2024b) performs data-free distillation on the server. To complete missing connections between subgraphs, FedSage (Zhang et al., 2021) and FedDEP (Zhang et al., 2024) additionally train a neighborhood generator. To better represent the minority in local graph data, FedSpray (Fu et al., 2024) learns class-wise structure proxies to mitigate biased neighborhoods. But it needs numerous communication rounds for optimization.
@@ 2.2 One-shot Federated Learning
One-shot federated learning largely reduces communication costs and circumvents potential man-in-the-middle attacks. Mainstream OFL methods can be classified into 3 categories. (1) Ensemble-based methods (Guha et al., 2019; Zhang et al., 2022; Dai et al., 2024), (2) Distillation-based methods (Zhou et al., 2020; Song et al., 2023), (3) Generative-based methods (Heinbaugh et al., 2023; Yang et al., 2024a, b). However, existing OFL methods primarily focus on image data and are either incompatible or ineffective for graph learning.
@@ 3.1 Notations and Definitions
Consider a federated graph learning setting with $m$ clients, where $k$-th client possesses the local undirected graph $G_{k}(\mathcal{V}_{k},\mathcal{E}_{k})$ with $|\mathcal{E}_{k}|$ edges and $|\mathcal{V}_{k}|=N_{k}$ nodes. Node features matrix is $\mathbf{X}_{k}=\{\mathbf{x}_{1},...,\mathbf{x}_{n_{k}}\}\in\mathbb{R}^{|\mathcal{V}_{k}|\times f}$, where $f$ denotes the feature dimension and $\mathbf{x}_{i}$ is the feature of node $v_{i}$. The adjacency matrix is $\mathbf{A}_{k}$ and augmented normailzied adjacency matrix is $\tilde{\mathbf{A}}_{k}$. The labels of nodes in this local graph are denoted by $\mathbf{Y}_{k}=\{y_{1},...,y_{n_{k}}\}\in\mathbb{R}^{|\mathcal{V}_{k}|\times C}$, where $C$ denotes the number of distinct classes. In the semi-supervised node classification task (Kipf & Welling, 2016), nodes on $k$-th client are partitioned into labeled nodes $\mathcal{V}_{k,L}$ and unlabeled nodes $\mathcal{V}_{k,U}$. The corresponding labels are $\mathbf{Y}_{k,L}$ and $\mathbf{Y}_{k,U}$, respectively. The neighborhood of node $v_{i}$ in $G_{k}(\mathcal{V}_{k},\mathcal{E}_{k})$ is defined as $\mathcal{N}_{v_{i}}=\{u|\exists e_{u,v_{i}}\in\mathcal{E}_{k}\}$, where edge $e_{u,v_{i}}$ connects node $u$ and $v_{i}$.
@@
Node Homophily. For a labeled node $v_{i}\in\mathcal{V}_{L}$ in an undirected graph $G(\mathcal{V},\mathcal{E})$, The node homophily $h_{node}(v_{i})$ quantifies the proportion of its labeled neighbors that share the same label $y_{v_{i}}$:
@@
$H(c)=\sum_{u\in\mathcal{N}_{c}^{G}}h_{node}(u),$ (2)
where $\mathcal{N}_{c}^{G}=\{v_{i}\in\mathcal{V}_{L}|y_{i}=c\}$ represents the set of labeled nodes in $G$ belonging to class $c$. Note that $H(c)$ considers both the quantity and topology of nodes. A higher $H(c)$ would indicate nodes of class $c$ are numerous and exhibit strong homophily, making them more predictable for a vanilla graph model (Kipf & Welling, 2016).
@@
which can be organized into a vector $\mathbf{w_{dist}}\in\mathbb{R}^{c}$ where the $i$-th entry corresponds to $w_{dist}(i)$ for class $i$. $\mathbf{w_{dist}}$ is used to determine $\gamma_{i}$ for each node, as described in Section 4.4.
@@ 3.2 Problem Formulation
Assume a situation where a whole graph is distributed on $m$ clients, each of which possesses its local non-overlapped subgraph. We formulate the one-shot personalized federated graph learning for node classification problem as:
@@ 4.2 Feature Distribution Estimation
$\boldsymbol{\mu}_{k}^{c}=$ $\frac{1}{|\mathcal{V}_{k,L}^{c}|}\sum_{v_{i}\in\mathcal{V}_{k,L}^{c}}\mathbf{x}^{prop}_{i},$ (6)
$\boldsymbol{s^{2}}_{k}^{c}=$ $\frac{1}{|\mathcal{V}_{k,L}^{c}|-1}\sum_{v_{i}\in\mathcal{V}_{k,L}^{c}}(\mathbf{x}^{prop}_{i}-\boldsymbol{\mu}_{k}^{c})^{2},$
@@
However, due to the limited number of labeled nodes, the estimation may lack accuracy. To address this, we propose a Homophily-guided Reliable node Expansion (HRE) strategy as a plugin to further augment the estimation process on homophilic graphs. In HRE, we first apply Label Propagation (Iscen et al., 2019) to obtain the soft labels $\tilde{\mathbf{y}}\in\mathbb{R}^{c}$ for unlabeled nodes. We denote $c_{i}^{\prime}=\arg\max_{0\leq i<c}\tilde{\mathbf{y}}(i)$, then identify reliable soft labels based on the following criteria: (C1) High Confidence: $\tilde{\mathbf{y}}_{i}(c_{i}^{\prime})\geq f_{th}$; (C2) High Class Homophily: $c_{i}^{\prime}\in topK(H)$, indicating the node is likely to share the same label with its neighbors; (C3) High Node Degree: $d(v_{i})\geq d_{th}$. A node with more neighbors could reduce the variance in predictions. With these criteria, the reliable nodes $\mathcal{V}_{k,r}$ set with inferred true labels is:
$\mathcal{V}_{k,r}=$$\{v_{i}|(v_{i}\in\mathcal{V}_{k,U})\wedge(\tilde{\mathbf{y}}_{i}(c^{\prime})\geq f_{th})\wedge$ (7)
$(d(v_{i})\geq d_{th})\wedge(c^{\prime}\in topK(H))\},$
@@
$\boldsymbol{\mu}_{k}^{c}=$ $\frac{1}{|\mathcal{V}_{k,L}^{c}\cup\mathcal{V}_{k,r}^{c}|}\sum_{v_{i}\in\mathcal{V}_{k,L}^{c}\cup\mathcal{V}_{k,r}^{c}}\mathbf{x}^{prop}_{i},$ (8)
$\boldsymbol{s^{2}}_{k}^{c}=$ $\frac{1}{|\mathcal{V}_{k,L}^{c}\cup\mathcal{V}_{k,r}^{c}|-1}\sum_{v_{i}\in\mathcal{V}_{k,L}^{c}\cup\mathcal{V}_{k,r}^{c}}(\mathbf{x}^{prop}_{i}-\boldsymbol{\mu}_{k}^{c})^{2},$

### v6

@@ Abstract
Federated Graph Learning (FGL) has emerged as a promising paradigm for breaking data silos among distributed private graphs. In practical scenarios involving heterogeneous distributed graph data, personalized Federated Graph Learning (pFGL) aims to enhance model utility by training personalized models tailored to client needs. However, existing pFGL methods often require numerous communication rounds under heterogeneous graphs, leading to significant communication overhead and security concerns. While One-shot Federated Learning (OFL) enables collaboration in a single round, existing OFL methods are designed for image-centric tasks and are ineffective for graph data, leaving a critical gap in the field. Additionally, personalized models derived from existing methods suffer from bias, failing to effectively generalize to the minority. To address these challenges, we propose the first One-shot personalized Federated Graph Learning method (O-pFGL) for node classification, compatible with Secure Aggregation protocols for privacy preservation. Specifically, for effective graph learning in one communication round, our method estimates and aggregates class-wise feature distribution statistics to construct a global pseudo-graph on the server, facilitating the training of a global graph model. To mitigate bias, we introduce a two-stage personalized training approach that adaptively balances local personal information and global insights from the pseudo-graph, improving both personalization and generalization. Extensive experiments on 12 multi-scale graph datasets demonstrate that our method significantly outperforms state-of-the-art baselines across various settings.
@@ 1 Introduction
Graphs are widely employed to model complex relationships between entities across a variety of domains [20, 25, 3]. While various algorithms and models [16, 26, 50] have been proposed, most of them assume that graph data from different sources are managed centrally. However, collecting and managing such graph data is often costly, impractical, and poses significant privacy risks in many sensitive scenarios [14].
@@
Figure 1: (a) A toy example explains the major classes and minor classes. (b) The empirical analysis of a client on the CiteSeer dataset under the Louvain partition to illustrate L3. Left: Existing methods result in a biased model that neglects the minority. Middle: The performance of existing methods and our method in class 1 with 50 samples. Right: The performance of existing methods and our method in class 2 with 14 samples. Our method effectively mitigates the performance gap.
@@
To address the tension between the need for vast datasets and the growing demand for privacy protection, Federated Graph Learning (FGL) has emerged as a promising solution, which enables collaborative training on distributed graph data without requiring centralized management [37, 17]. While FGL shows its potential, it faces two major limitations: (1) Typical FGL struggles to effectively personalize models for joint improvement across heterogeneous clients’ graph data [2]. This heterogeneity arises from variations of both node and structural properties, with adjacent nodes influencing each other, further complicating personalization. (2) FGL requires iterative communication between clients and the server, raising efficiency concerns due to communication overhead [28, 8] and security concerns (e.g. man-in-the-middle attacks [47, 36, 67]). These limitations can be mitigated by integrating personalized federated graph learning (pFGL) and one-shot federated learning (OFL), which offer novel solutions to improve communication efficiency and model personalization.
@@
To address these challenges, we explore the integration of pFGL and OFL. The pFGL methods allow for personalized model training by focusing on individual client graph data. OFL methods limit communication to a single round, thereby reducing the communication overhead and circumventing security risks such as man-in-the-middle attacks [18, 62]. However, both existing pFGL and OFL methods face challenges in FGL. Specifically, they exhibit the following limitations:
@@
L1: Ineffectiveness for graph learning in one communication round. Existing pFGL methods [2, 33, 69] rely on iterative communication for optimization, making them ineffective within one communication round. Most methods also fail to support heterogeneous models, as they primarily rely on aggregating model parameters. Meanwhile, existing OFL methods are incompatible with graph data since they focus primarily on image-centric tasks. The ensemble-based OFL methods [62, 9] can generate images from ensemble models, but struggle with graph data due to the inter-dependencies between nodes. The distillation-based OFL faces the problems of distilling inter-related graphs. The generative-based OFL methods [21, 55, 54] rely on pre-trained generative models (e.g., Stable Diffusion [38]), which are impractical for graph generation.
@@
L2: Incompatibility with Secure Aggregation. Secure Aggregation is a widely used technique [4, 41] designed to safeguard client privacy during the federated aggregation process. However, most pFGL and OFL methods require clients to upload raw model parameters or auxiliary information independently, which complicates the aggregation process (e.g., ensemble or similarity calculation) and hinders the implementation of privacy-preserving weighted averages.
@@
L3: Imbalance between personalization and local generalization. Local graph data often exhibit a significant imbalance in both quantity and topology, as illustrated in Figure 1. Personalized models tend to be biased toward major classes (with larger quantities and high homophily to form compact communities), while neglecting minor classes with fewer samples and often surrounded by nodes from major classes. This imbalance leads to a degradation in model performance, particularly in underrepresented classes. A detailed empirical analysis is provided in Figure 1, which shows that existing personalized methods result in heavy bias toward majority classes (class 3 and class 5, with a total of 180 samples) while failing to adequately represent minority classes (class 1 and class 2, with only 64 samples), highlighting the limitations of current methods in handling imbalanced graph data.
@@
To overcome these limitations, we propose a One-shot personalized Federated Graph Learning method (O-pFGL). Our method supports model heterogeneity and Secure Aggregation protocols for privacy preservation. Specifically, for L1 and L2, clients upload estimated feature distribution statistics, augmented by our proposed homophily-guided reliable node expansion strategy, instead of raw model parameters. The server securely aggregates these statistics to generate a global pseudo-graph, which is subsequently utilized to train personalized models on clients, eliminating the need for direct model parameter exchanges and preventing model theft. To address L3, we propose a two-stage personalized training approach. In the first stage, clients train a global generalized model based on the pseudo-graph, which serves as both a fixed teacher capturing global knowledge and a start-point for subsequent fine-tuning. In the second stage, clients fine-tune the global model on local graph data with the node-adaptive distillation, which leverages the global knowledge from the teacher model to counteract biases toward the majority, ensuring effective learning for the minority and achieving a robust balance between personalization and generalization.
@@
Our method significantly outperforms state-of-the-art baselines through comprehensive experiments across various settings, including different graph scales, partition strategies, learning paradigms, and model architectures. Experimental results demonstrate the superiority of our method in achieving a one-shot personalized federated graph learning solution that balances efficiency, privacy, and personalization. To sum up, our contributions are listed as follows:
@@ 2.1 Federated Graph Learning
From the graph level [52, 43], each client possesses multiple independent graphs (e.g., molecular graphs) and targets at graph-level tasks. From the subgraph level, each client possesses a subgraph of an implicit global graph. FedPUB [2] and FedGTA [33] personalize models by similarity of functional embedding and mixed moments of neighbor features. AdaFGL [32] studies the structure non-IID problem. FedTAD [69] performs data-free distillation on the server. To complete missing connections between subgraphs, FedSage [63] and FedDEP [64] additionally train a neighborhood generator. To better represent the minority in local graph data, FedSpray [15] learns class-wise structure proxies to mitigate biased neighborhoods. But it needs numerous communication rounds for optimization. FedLoG [30] synthesizes global data from clients’ synthetic nodes and generalizes the local training. Detailed description and critical analysis are in the Appendix A.6.
@@ 2.2 One-shot Federated Learning
One-shot federated learning largely reduces communication costs and circumvents potential man-in-the-middle attacks. Mainstream OFL methods can be classified into: (1) Ensemble-based methods [18, 62, 9], (2) Distillation-based methods [66, 42], (3) Generative-based methods [21, 54, 55]. However, existing OFL methods primarily focus on image data and are either incompatible or ineffective for graph learning. Detailed description and critical analysis are in the Appendix A.6.
@@ 3.1 Notations and Definitions
Consider a federated graph learning setting with $m$ clients, where $k$-th client possesses the local undirected graph $G_{k}(\mathcal{V}_{k},\mathcal{E}_{k})$ with $|\mathcal{E}_{k}|$ edges and $|\mathcal{V}_{k}|=N_{k}$ nodes. Node features matrix is $\mathbf{X}_{k}=\{\mathbf{x}_{1},...,\mathbf{x}_{n_{k}}\}\in\mathbb{R}^{|\mathcal{V}_{k}|\times f}$, where $f$ is the feature dimension and $\mathbf{x}_{i}$ is the feature of node $v_{i}$. The adjacency matrix is $\mathbf{A}_{k}$ and the augmented normalized adjacency matrix is $\tilde{\mathbf{A}}_{k}$. The labels of nodes in its local graph are denoted by $\mathbf{Y}_{k}=\{y_{1},...,y_{n_{k}}\}\in\mathbb{R}^{|\mathcal{V}_{k}|\times C}$, where $C$ is the number of distinct classes. In the semi-supervised node classification task [31], nodes on the $k$-th client are partitioned into labeled nodes $\mathcal{V}_{k,L}$ (i.e., nodes in the training set) and unlabeled nodes $\mathcal{V}_{k,U}$ (i.e., their labels are unknown during training). The corresponding labels are $\mathbf{Y}_{k,L}$ and $\mathbf{Y}_{k,U}$, respectively. The neighborhood of node $v_{i}$ in $G_{k}(\mathcal{V}_{k},\mathcal{E}_{k})$ is defined as $\mathcal{N}_{v_{i}}=\{u|\exists e_{u,v_{i}}\in\mathcal{E}_{k}\}$, where edge $e_{u,v_{i}}$ connects node $u$ and $v_{i}$.
@@
Node Homophily. For a labeled node $v_{i}\in\mathcal{V}_{L}$ in an undirected graph $G(\mathcal{V},\mathcal{E})$, the node homophily $h_{node}(v_{i})$ quantifies the proportion of its labeled neighbors that share the same label:
@@
$H(c)=\sum_{u\in\mathcal{N}_{c}}h_{node}(u),$ (2)
where $\mathcal{N}_{c}=\{v_{i}\in\mathcal{V}_{L}|y_{i}=c\}$ is the set of labeled nodes belonging to class $c$. Note that $H(c)$ considers both the quantity and topology of nodes. A higher $H(c)$ indicates nodes of class $c$ are numerous and exhibit strong homophily, making them more predictable for a vanilla graph model [31].
@@
which can be organized into a vector $\mathbf{w}_{dist}\in\mathbb{R}^{C}$ where the $i$-th entry corresponds to $w_{dist}(i)$ for class $i$. $\mathbf{w}_{dist}$ is used to determine node-adaptive distillation factor $\gamma_{i}$, as described in Section 4.4.
@@ 3.2 Problem Formulation
Assume a situation where a whole graph is distributed on $m$ clients, each of which possesses its local non-overlapped subgraph. We formulate the one-shot personalized federated graph learning for the node classification problem as:
@@ 4.2 Feature Distribution Estimation
$\boldsymbol{\mu}_{k}^{c}=\frac{1}{|\mathcal{V}_{k,L}^{c}|}\sum_{v_{i}\in\mathcal{V}_{k,L}^{c}}\mathbf{x}^{prop}_{i},\quad\boldsymbol{s^{2}}_{k}^{c}=\frac{1}{|\mathcal{V}_{k,L}^{c}|-1}\sum_{v_{i}\in\mathcal{V}_{k,L}^{c}}(\mathbf{x}^{prop}_{i}-\boldsymbol{\mu}_{k}^{c})^{2},$ (6)
@@
However, the estimation may lack accuracy due to the limited number of labeled nodes. To address this, we propose a Homophily-guided Reliable node Expansion (HRE) strategy as a plugin to further augment the estimation process on homophilic graphs. In HRE, we first apply Label Propagation [26] to obtain the soft labels $\tilde{\mathbf{y}}\in\mathbb{R}^{C}$ for unlabeled nodes. We denote $c_{i}^{\prime}=\arg\max_{0\leq i<C}\tilde{\mathbf{y}}(i)$, then identify reliable soft labels based on the following criteria: (C1) High Node Degree: $d(v_{i})\geq d_{th}$; (C2) High Confidence: $\tilde{\mathbf{y}}_{i}(c_{i}^{\prime})\geq f_{th}$; (C3) High Class Homophily: $c_{i}^{\prime}\in topK(H)$, indicating the node is likely to share the same label with its neighbors. A node with more neighbors could reduce the variance in predictions. With these criteria, the reliable nodes $\mathcal{V}_{k,r}$ set with inferred true labels is:
$\mathcal{V}_{k,r}=\{v_{i}|(v_{i}\in\mathcal{V}_{k,U})\wedge(d(v_{i})\geq d_{th})\wedge(\tilde{\mathbf{y}}_{i}(c^{\prime})\geq f_{th})\wedge(c^{\prime}\in topK(H))\},$ (7)
@@
$\boldsymbol{\mu}_{k}^{c}=\frac{1}{|\mathcal{V}_{k,L}^{c}\cup\mathcal{V}_{k,r}^{c}|}\sum_{v_{i}\in\mathcal{V}_{k,L}^{c}\cup\mathcal{V}_{k,r}^{c}}\mathbf{x}^{prop}_{i},\quad\boldsymbol{s^{2}}_{k}^{c}=\frac{1}{|\mathcal{V}_{k,L}^{c}\cup\mathcal{V}_{k,r}^{c}|-1}\sum_{v_{i}\in\mathcal{V}_{k,L}^{c}\cup\mathcal{V}_{k,r}^{c}}(\mathbf{x}^{prop}_{i}-\boldsymbol{\mu}_{k}^{c})^{2},$ (8)